# Simulating and probing many-body quantum states in waveguide-QED systems with giant atoms

### v1

@@ III.1 A scheme for constructing an AAH-type atomic chain using decoherence-free interactions
Now we construct the AAH-type atomic chain based on the decoherence-free interactions between each pair of braided atoms, which is a unique property in giant-atom systems Kockum et al. (2018). To this end, we set $\gamma_{1}=\gamma_{2}=\gamma$ (i.e., $\delta=0$, together with $\theta_{i2}-\theta_{i1}=\pi$, the condition for decoherence-free interactions between nearest neighbor atoms is satisfied). From Eqs. (11) and (12b), it can be seen that the individual decays and the collective decays all vanish, with $\Gamma_{\mathrm{eff}}=\Gamma^{(\mathrm{coll})}_{ij}=0$. And for exchange interactions, only the nearest-neighbor ones are nonzero, with $G_{ij}=\gamma\sin\phi\equiv J$ for $|i-j|=1$, and $G_{ij}=0$ for $|i-j|>1$ [see Eq. (12a)]. Clearly, under these parameters the effective Hamiltonian (10) becomes the standard diagonal AAH Hamiltonian
$\hat{H}_{\mathrm{AAH}}$ $=$ $V_{0}\sum_{i=1}^{N}\cos\left(2\pi\beta{i}+\varphi\right)\hat{\sigma}^{+}_{i}\hat{\sigma}^{-}_{i}$ (13)
$+J\sum_{i=1}^{N-1}\left(\hat{\sigma}^{+}_{i}\hat{\sigma}^{-}_{i+1}+\mathrm{H.c.}\right).$
Although a braided atomic chain under the above parameters can be exactly described by a diagonal AAH model, its state cannot be probed by the photon scattering spectra because all the atoms are decoupled from the waveguide. Thus the accuracy of the simulated model must be balanced against the detectability of the atomic chain. This necessitates that the decay rates of the two connection points within each atom should be slightly different, with $\delta\ll\gamma$. Under this condition, each atom can obtain a tiny effective decay, which can be approximated as
@@
Thus the $i$th atom can interact with the photonic modes with a strength $\mathcal{V}_{i}\sim{\sqrt{{\Gamma_{\mathrm{eff}}}/{2}}}$. Furthermore, it can be readily demonstrated that for $\delta\ll\gamma$, the atom array can still be approximately described by the AAH model. Without loss of generality, we set the phase factor to be $\phi=\pi/2$, so the nearest-neighbor interaction strength become $J=\gamma$. In addition, for non-nearest-neighbor interactions (with $|i-j|>1$), we have $|G_{ij}|/J=\Gamma_{\mathrm{eff}}/(2\gamma)\simeq\delta^{2}/(2\gamma^{2})$ for $|i-j|\in\mathbb{O}^{+}$ and $G_{ij}=0$ for $|i-j|\in\mathbb{E}^{+}$ [see Eqs. (12a)]. For collective decay rates, we have $\Gamma^{(\mathrm{coll})}_{ij}=0$ for $|i-j|\in\mathbb{O}^{+}$ and $|\Gamma^{(\mathrm{coll})}_{ij}|/J=\Gamma_{\mathrm{eff}}/\gamma\simeq\delta^{2}/\gamma^{2}$ for $|i-j|\in\mathbb{E}^{+}$ [see Eqs. (12b)]. These results demonstrate that, in comparison to nearest-neighbor coherent interactions, the non-nearest-neighbor coherent interactions and all dissipative interactions are either negligible or absent. Thus the atomic chain can still be approximately described by the AAH Hamiltonian (13), but each atom weakly couples to the waveguide modes. This implies that the braided atomic chain sketched in Fig. 1(a) is equivalent to the configuration shown in Fig. 1(b). The corresponding scattering amplitudes for incident photons can be calculated by using Eqs. (6a) and (6b).
@@
(i) The interactions between atoms are mediated by the waveguide modes, which are realized by designing the layout of the connection points. In contrast to conventional methods, direct couplings between atoms are not required.
(ii) In addition to inducing atom-atom interactions, the waveguide also acts as a measurement device, enabling the probing of the energy spectra and quantum states of the atomic chain through the photon scattering process.
(iii) In conventional spectroscopic technique Roushan et al. (2017); Shi et al. (2023) for detecting the spectral structure of a qubit array, it is necessary to perform the measurement on each atom in sequence. This is because measuring a single atom only provides partial information about the energy spectrum. In contrast, our method enables the complete energy spectrum to be revealed by a single scattering process, as all atoms are coupled to the waveguide. Thus, our proposed measurement scheme may prove to be a highly efficient alternative to existing methodologies.
III.2 Probing the energy band structure through single-photon scattering spectra
@@
Here $\tilde{\Delta}_{n}=\tilde{\omega}_{n}-\omega_{\mathrm{a}}$ is the detuning between the $n$th collective mode and the atomic frequency. $\tilde{\Gamma}_{n}$ denotes the effective decay of the $n$th collective mode. ${\eta}_{n}$ determines the weight of the $n$th Lorentzian component. The obtained Lorentzian components are illustrated in the lower panel of Fig. 3(a).
It can be verified that the center frequencies of the Lorentzian components fit well with the energy levels given by the exact AAH model [see Fig. 3(b)]. Additionally, as illustrated in Fig. 3(c), only one of the bulk states in each band is superradiant (with $\tilde{\Gamma}_{n}>\Gamma_{\mathrm{eff}}$), while the remainder are subradiant (with $\tilde{\Gamma}_{n}<\Gamma_{\mathrm{eff}}$). The superposition and interference between these Lorentzian amplitudes give rise to the transmission spectrum depicted in the upper panel of Fig. 3(a). Finally, the decay rate of the two edge states in the gaps is approximately equal to the effective decay rate $\Gamma_{\mathrm{eff}}$ of a single atom [see Fig. 3(c)], given that in these states, only the atom at one end is highly excited. It should be noted that the assumptions of $\delta\ll\gamma$ (resulting in $\Gamma_{\mathrm{eff}}\simeq\delta^{2}/\gamma$) and $\phi=\pi/2$ (resulting in $J=\gamma$) have been made to ensure that the atom array can be approximately described by the AAH model. Consequently, the relation $N\Gamma_{\mathrm{eff}}/J\simeq N\delta^{2}/\gamma^{2}\ll 1$ can be easily satisfied for not so large $N$. Namely, the widths of the resonances (the majority of which are smaller than the value of $\Gamma_{\mathrm{eff}}$) are considerably less than the spectrum width (which is of the order of $J$) divided by $N$. This property offers a high resolution for the transmission dips associated with the energy levels of the atomic chain, thereby facilitating precise probing of the band structures, as shown in Fig. 3(a). The following section will further illustrate the potential for simulating the Hofstadter butterfly spectrum by exploiting this advantage of giant-atom wQED.
@@ III.3 Reconstructing the fractal structure of the Hofstadter butterfly spectrum
It is well known that the problem of Bloch electrons on a 2D lattice subject to a perpendicularly applied magnetic field is described by the Hofstadter model. When the magnetic field reaches a strength that allows one magnetic flux quantum per unit cell to pass through, the corresponding energy spectrum versus the dimensionless magnetic flux form a famous fractal structure known as the Hofstadter butterfly Hofstadter (1976). Nevertheless, a magnetic field of approximately $10^{4}$ is necessary to observe the Hofstadter butterfly in typical crystals, which is beyond the capabilities of current experimental techniques. One potential solution to this challenge is the use of superlattices with a larger lattice constant, which could reduce the requisite strength of magnetic fields Dean et al. (2013); Ponomarenko et al. (2013); Hunt et al. (2013). On the other hand, the 2D Hofstadter model can be exactly mapped to the 1D diagonal AAH model. Consequently, the Hofstadter butterfly spectrum has also been successfully simulated in 1D AAH-type chains based on superconducting circuits Shi et al. (2023); Roushan et al. (2017). In this approach, the modulation frequency $\beta$ is used to analogize the required magnetic flux, thereby eliminating the need for a complex 2D lattice and an external magnetic field.
@@
As previously outlined in Sec. III.1, in our proposal on generating a diagonal AAH model based on giant-atom wQED, the waveguide modes serve two distinct purposes. Firstly, they mediate decoherence-free interactions between neighboring atoms. Secondly, they provide channels for probing the energy spectrum of the atom array. Moreover, as demonstrated in Sec. III.2, the majority of the resonance dips corresponding to the energy levels of the AAH model can be well resolved in the transmission spectrum. Therefore, the scattering spectrum of this system can be effectively utilized to reconstruct the complex and fine fractal structure of the Hofstadter butterfly spectrum with a high degree of fidelity, as shown in Fig. 4(a). In this panel, the number of giant atoms is set to $N=30$, and the reduced magnetic flux $\beta$ is varied from 00 to $1$. This generates $299$ distinct instances of the reflection spectrum, which exhibit clear fractal features and a shape that is reminiscent of a butterfly. Figure 4(b) depicts the Hofstadter butterfly energy spectrum obtained from the theoretical model, which exhibits a high degree of agreement with the result in Fig. 4(a). Thus, it appears that an alternative or potentially more efficient approach to simulating the Hofstadter butterfly spectrum is feasible.
Figure 5: (a) The square of the modulus of the wave functions of AAH model with $N=30$, $V_{0}=J/2$, $\beta=(\sqrt{5}-1)/2$, and $\varphi=0$. (b) The transmission spectrum (upper panel) and corresponding Lorentzian decompositions (lower panel) for a wQED system with an AAH-type giant-atom array. The phase delay is selected to be $\phi=\pi/2$, resulting in $J=\gamma$. The quantity $\delta$ is chosen as $\delta=0.1\gamma$. The size of the array and the on-site modulations are the same as in (a). In the lower panel, the resonances corresponding to the superradiant (subradiant) states are labeled as SupS (SubS). (c) and (d) Similar to (a) and (b), but with $V_{0}=8J$. (e) snd (f) The decay rate $\tilde{\Gamma}_{n}$ of each Lorentzian component for $V_{0}=J/2$ and $V_{0}=8J$. The effective decay $\Gamma_{\mathrm{eff}}$ of a single giant atom is labeled by the dashed line.
Figure 6: $\sigma^{2}$ and $\mathrm{IPR}_{\mathrm{decay}}$ as functions of $V_{0}$, with (a) $N=30$, $\delta=0.1\gamma$, and (b) $N=300$, $\delta=0.01\gamma$. Other parameters are the same as those used in Fig. 5. The extended (E) and the localized (L) phases are well characterized by these quantities. In panels (a) and (b), the inverse participation ratio $\mathrm{IPR}(1)$ (Here 1 is used to denote the ground state) of the standard AAH chain is also provided for comparison.
@@ III.4 Probing the localization transition
The localization transition Aubry and André (1980); Simon (1982); Jitomirskaya (1999) is another interesting topic in an incommensurate diagonal AAH model. When the on-site potential of the diagonalized AAH model is made quasiperiodic by setting $\beta$ to an irrational number, the model exhibits a localization transition, where all bulk eigenstates are extended for $V_{0}<2J$, and localized for $V_{0}>2J$. Here we will present an intriguing correlation between the localization transition of the atomic chain and the photon transport behavior observed in the waveguide.
@@
To analyze the relation between the localization transition and the scattering spectrum, we let the value of modulation frequency irrational, with $\beta=(\sqrt{5}-1)/2$, and present in Fig. 5(a) the square of the modulus of the wave functions $\psi_{m}^{(n)}$ (where $m$ and $n$ are used to index the sites and eigenstates, respectively) of the diagonal AAH model for $V_{0}=J/2$, which corresponds to the extended phase. The resulting transmission spectrum and the corresponding Lorentzian decompositions are illustrated in Fig. 5(b). These collective modes with extended feature comprise a small number of superradiant states (with $\tilde{\Gamma}_{n}>\Gamma_{\mathrm{eff}}$) and a significant proportion of subradiant states (with $\tilde{\Gamma}_{n}<\Gamma_{\mathrm{eff}}$), exhibiting a wide range of linewidths, as illustrated in the lower panel of Fig. 5(b). While for $V_{0}=8J$, the situation is quite different. As shown in Fig. 5(c), the wave functions of the atomic chain are highly localized. In this case, the decay of each collective mode to the waveguide is primarily attributable to the most excited atom, which has an excitation probability close to one. Therefore, the width of each resonance dip in the reflection spectrum [see the upper panel of Fig. 5(d)] and the width of each collective excitation amplitude [see the lower panel of Fig. 5(d)] are approximately equal to the effective decay rate $\Gamma_{\mathrm{eff}}$ of a single atom. Namely, all collective states exhibit neither a markedly superradiant nor a distinctly subradiant character.
@@
A more straightforward illustration of the decay rates of the collective modes as a function of index $n$ in the extended and localized phases can be found in Figs. 5(e) and 5(f), respectively. It can be observed that there exists an interesting correspondence between the extended (localized) phase of the atom array and the “localization” (“delocalization”) of the decay rates of its collective modes. We now seek ways to quantitatively characterize this correspondence. Note that the degree of localization of the $n$th eigenstate of the AAH chain can be described by the inverse participation ratio (IPR):
$\mathrm{IPR}(n)=\sum_{m}\left|\psi^{(n)}_{m}\right|^{4},$ (16)
@@
which vanishes for extended states [see the curves marked with squares in Figs. 6(a) and 6(b)]. As for the aspect of the collective modes, the mean value of their decay rates is equal to the decay rate of a single atom, with $(\sum_{n}\tilde{\Gamma}_{n})/N=\Gamma_{\mathrm{eff}}$. It is therefore reasonable to conclude that the variance
$\sigma^{2}=\frac{1}{N}\sum_{n}\left(\tilde{\Gamma}_{n}-\Gamma_{\mathrm{eff}}\right)^{2}$ (17)
@@
may be employed as a means of measuring the concentration of the dissipation-rate distribution (vs index $n$). It is evident that the variance $\sigma^{2}\simeq 0$ when $V_{0}>2J$, and $\sigma^{2}>0$ when $V_{0}<2J$ [see the curves marked with circles in Fig. 6(a) and 6(b)], indicating the localized and extended phases, respectively. Note that when $V_{0}\to 0$, the order of magnitude of $\sigma^{2}$ is $N\Gamma_{\mathrm{eff}}^{2}$ for an array with $N\gg 1$. This is due to the fact that in this regime, the decay rates of the superradiant modes, which are a few in number and of the order of $N\Gamma_{\mathrm{eff}}$, contribute mainly to the variance. One can further define a “normalized” decay rate $\tilde{\Gamma}^{\prime}_{n}=\tilde{\Gamma}_{n}/(N\Gamma_{\mathrm{eff}})$, with $\sum_{n}\tilde{\Gamma}^{\prime}_{n}=1$ (which is analogous to the normalization condition $\sum_{m}|\psi^{(n)}_{m}|^{2}=1$ of the wave functions). Therefore, it is possible to emulate the IPR for the wave functions [see Eq. (16)] to define the following IPR describing the distribution of the decay rates
$\mathrm{IPR}_{\mathrm{decay}}=\sum_{n}\tilde{\Gamma}^{\prime 2}_{n}.$ (18)
@@
The inverse of this quantity indicates the number of the collective modes that can decay appreciably into the waveguide. As shown in Fig. 6, both $\sigma^{2}$ and $\mathrm{IPR}_{\mathrm{decay}}$ can be used to quantify the localization transition. Moreover, in the thermodynamic limit $N\to\infty$, we have $\sigma^{2}=N\Gamma^{2}_{\mathrm{eff}}\times\mathrm{IPR}_{\mathrm{decay}}$. This is consistent with the results (see the curves marked with circles and triangles) in Fig. 6(b), where a relatively large value of $N=300$ is chosen. Besides, a comparison of Figs. 6(a) and 6(b) illustrates that a finite-size system comprising a greater number of sites is more effective at exhibiting characteristics associated with quantum phase transitions.

### v2

@@ III.1 A scheme for constructing an AAH-type atomic chain using decoherence-free interactions
Now we construct the AAH-type atomic chain based on the decoherence-free interactions between each pair of braided atoms, which is a unique property in giant-atom systems [35]. To this end, we set $\gamma_{1}=\gamma_{2}=\gamma$ (i.e., $\delta=0$, together with $\theta_{i2}-\theta_{i1}=\phi+\phi^{\prime}=(2n-1)\pi$, the condition for decoherence-free interactions between nearest neighbor atoms is satisfied). From Eqs. (11) and (12b), it can be seen that the individual decays and the collective decays all vanish, with $\Gamma_{\text{eff}}=\Gamma^{(\text{coll})}_{ij}=0$. And for exchange interactions, only the nearest-neighbor ones are nonzero, with $G_{ij}=\gamma\sin\phi\equiv J$ for $|i-j|=1$, and $G_{ij}=0$ for $|i-j|>1$ [see Eq. (12a)]. Clearly, under these parameters the effective Hamiltonian (10) becomes the standard diagonal AAH Hamiltonian
$\hat{H}_{\text{AAH}}$ $=$ $V_{0}\sum_{i=1}^{N}\cos\left(2\pi\beta{i}+\varphi\right)\hat{\sigma}^{+}_{i}\hat{\sigma}^{-}_{i}$ (13)
@@
It should be emphasized that the phase delays $\phi$ and $\phi^{\prime}$ can be manipulated by modulating the transition frequencies of the artificial atoms (e.g., through flux lines coupled to qubits in superconducting circuits [6]). This manipulation can produce a range of values for the waveguide-induced decoherence-free interaction. Note that in experiments, the ratios $\zeta=\phi^{\prime}/\phi$ are fixed in hardware by the relative lengths of the waveguide segments. Therefore, the condition $\phi+\phi^{\prime}=(2n-1)\pi$ for decoherence-free interaction, under which the desired tight-binding atomic chain can be constructed, requires that the phase factor be tuned to $\phi=(2n-1)\pi/(\zeta+1)$, resulting in
$J=\gamma\sin\left[\frac{(2n-1)\pi}{\zeta+1}\right].$ (14)
This allows us to get multiple values for the coupling strength $J\in[-\gamma,\gamma]$ (especially for relatively large $\zeta$), making the spectral width of the 1D chain highly tunable. Without loss of generality, in the following discussion we set the phase factor to be $\phi=\phi^{\prime}=\pi/2$ (i.e., $\zeta=1,n=1$), so that the nearest-neighbor interaction strength reaches its maximum value $J=\gamma$.
Although a braided atomic chain under the above parameters can be exactly described by the diagonal AAH model, its states cannot be probed by photon scattering spectrum because all the atoms are decoupled from the waveguide. Thus the accuracy of the simulated model must be balanced against the detectability of the atomic chain. This necessitates that the decay rates of the two connection points within each atom should be slightly different, with $\delta\ll\gamma$. Under this condition, each atom can obtain a tiny effective decay, which can be approximated as
$\Gamma_{\text{eff}}\simeq\frac{\delta^{2}}{\gamma}\end{aligned}.$ (15)
Thus the $i$th atom can interact with the photonic modes with a strength $\mathcal{V}_{i}\sim{\sqrt{{\Gamma_{\text{eff}}}/{2}}}$. Furthermore, it can be readily demonstrated that for $\delta\ll\gamma$, the atom array can still be approximately described by the AAH model. First for $\phi=\pi/2$ and $\delta\ll\gamma$, the nearest-neighbor interaction strength is still $J=\gamma$ [see Eq. (12a)]. In addition, for non-nearest-neighbor interactions (with $|i-j|>1$), we have $|G_{ij}|/J=\Gamma_{\text{eff}}/(2\gamma)\simeq\delta^{2}/(2\gamma^{2})$ for $|i-j|\in\mathbb{O}^{+}$ and $G_{ij}=0$ for $|i-j|\in\mathbb{E}^{+}$ [see Eq. (12a)]. For collective decay rates, we have $\Gamma^{(\text{coll})}_{ij}=0$ for $|i-j|\in\mathbb{O}^{+}$ and $|\Gamma^{(\text{coll})}_{ij}|/J=\Gamma_{\text{eff}}/\gamma\simeq\delta^{2}/\gamma^{2}$ for $|i-j|\in\mathbb{E}^{+}$ [see Eq. (12b)]. These results demonstrate that, in comparison to nearest-neighbor coherent interactions, the non-nearest-neighbor coherent interactions and all dissipative interactions are either negligible or absent. Thus the atomic chain can still be approximately described by the AAH Hamiltonian (13), but each atom weakly couples to the waveguide modes. This implies that the braided atomic chain sketched in Fig. 1(a) is equivalent to the configuration shown in Fig. 1(b). The corresponding scattering amplitudes for incident photons can be calculated by using Eqs. (6a) and (6b).
Note that the above analysis is based on the assumption that the phase accumulation effects for detuned photons (the non-Markovian effects) are negligible. If these effects are included, the phase factor $\phi$ should be replaced by $(1+\Delta/\omega_{\text{a}})\phi$ [28]. Thus, for $\phi=\pi/2$ considered here, the accumulated phase factor between the leftmost and the rightmost connection points of the atomic chain is $(N+1)(1+\Delta/\omega_{\text{a}})\pi/2\simeq N(1+\Delta/\omega_{\text{a}})\pi/2$. Moreover, the detuning range we are interested in is approximately $\Delta\sim\gamma$. Therefore, the extra phase $N\Delta\pi/(2\omega_{\text{a}})$ due to the effects of detuning can be safely neglected if the condition
$N\gamma\ll\omega_{\text{a}}$ (16)
is satisfied. It is evident that for a 1D chain of tens to hundreds of atoms, this condition can be readily fulfilled within the currently available parameters of wQED systems [3, 6] (with $\omega_{\text{a}}/\gamma\sim 10^{3}-10^{4}$).
@@
(i) The interactions between atoms are mediated by the waveguide modes, thereby obviating the necessity for direct couplings between atoms. In wQED with giant atoms, this kind of phase-dependent interactions are easy to modulate [6, 35]. For instance, in our proposal, the tunability of the nearest-neighbor coupling strength for an AAH type tight-binding atomic chain can easily be implemented via a control line (e.g., a flux line [6]) to modulate the atomic frequency. However, for small atoms, the creation of a tight-binding atomic chain with analogous tunability can only be achieved through the use of tunable coupling elements (usually in the form of superconducting qubits or superconducting quantum interference devices [67, 68, 69]), which requires more complex circuits.
(ii) In addition to inducing atom-atom interactions, the waveguide also acts as a measurement device, enabling the probing of the energy spectra and quantum states of the atomic chain through the photon scattering process. Thus additional readout devices coupled to the artificial atoms (e.g., resonators in superconducting circuits) are not required. This is particularly advantageous for atomic chains of large size, as it leads to substantial savings in hardware resources.
(iii) In conventional spectroscopic technique [61, 62] for detecting the spectral structure of a qubit array, it is necessary to perform the measurement on each atom in sequence. This is because measuring a single atom only provides partial information about the energy spectrum. In contrast, our method enables the complete energy spectrum to be revealed by a single scattering process, as all atoms are coupled to the waveguide. Thus, our proposed measurement scheme may prove to be a highly efficient alternative to existing methodologies.
III.2 Probing energy band structure through single-photon scattering spectra
@@
Here $\tilde{\Delta}_{n}=\tilde{\omega}_{n}-\omega_{\text{a}}$ is the detuning between the $n$th collective mode and the atomic frequency. $\tilde{\Gamma}_{n}$ denotes the effective decay of the $n$th collective mode. ${\eta}_{n}$ determines the weight of the $n$th Lorentzian component. The obtained Lorentzian components are illustrated in the lower panel of Fig. 3(a).
It can be verified that the center frequencies of the Lorentzian components fit well with the energy levels given by the exact AAH model [see Fig. 3(b)]. Additionally, as illustrated in Fig. 3(c), only one of the bulk states in each band is superradiant (with $\tilde{\Gamma}_{n}>\Gamma_{\text{eff}}$), while the remainder are subradiant (with $\tilde{\Gamma}_{n}<\Gamma_{\text{eff}}$). The superposition and interference between these Lorentzian amplitudes give rise to the transmission spectrum depicted in the upper panel of Fig. 3(a). Finally, the decay rate of the two edge states in the gaps is approximately equal to the effective decay rate $\Gamma_{\text{eff}}$ of a single atom [see Fig. 3(c)], given that in these states, only the atom at one end is highly excited. It should be noted that the assumptions of $\delta\ll\gamma$ (resulting in $\Gamma_{\text{eff}}\simeq\delta^{2}/\gamma$) and $\phi=\pi/2$ (resulting in $J=\gamma$) have been made to ensure that the atom array can be approximately described by the AAH model. Consequently, the relation $N\Gamma_{\text{eff}}/J\simeq N\delta^{2}/\gamma^{2}\ll 1$ can be easily satisfied for not so large $N$. Namely, the widths of the resonances (the majority of which are smaller than the value of $\Gamma_{\text{eff}}$) are considerably less than the spectrum width (which is of the order of $J$) divided by $N$. This property offers a high resolution for the transmission dips associated with the energy levels of the atomic chain, thereby facilitating precise probing of the band structures, as shown in Fig. 3(a). The following section will further illustrate the potential for simulating the Hofstadter butterfly spectrum by exploiting this advantage of giant-atom wQED.
@@
Figure 5: (a) The square of the modulus of the wave functions of AAH model with $N=30$, $V_{0}=J/2$, $\beta=(\sqrt{5}-1)/2$, and $\varphi=0$. (b) The transmission spectrum (upper panel) and corresponding Lorentzian decompositions (lower panel) for a wQED system with an AAH-type giant-atom array. The phase delay is selected to be $\phi=\pi/2$, resulting in $J=\gamma$. The quantity $\delta$ is chosen as $\delta=0.1\gamma$. The size of the array and the on-site modulations are the same as in (a). In the lower panel, the resonances corresponding to the superradiant (subradiant) states are labeled as SupS (SubS). (c) and (d) Similar to (a) and (b), but with $V_{0}=8J$. (e) snd (f) The decay rate $\tilde{\Gamma}_{n}$ of each Lorentzian component for $V_{0}=J/2$ and $V_{0}=8J$. The effective decay $\Gamma_{\text{eff}}$ of a single giant atom is labeled by the dashed line.
Figure 6: $\sigma^{2}$ and $\text{IPR}_{\text{decay}}$ as functions of $V_{0}$, with (a) $N=30$, $\delta=0.1\gamma$, and (b) $N=300$, $\delta=0.01\gamma$. Other parameters are the same as those used in Fig. 5. The extended (E) and the localized (L) phases are well characterized by these quantities. In panels (a) and (b), the inverse participation ratio $\text{IPR}(1)$ (Here 1 is used to denote the ground state) of the standard AAH chain is also provided for comparison.
@@ III.3 Reconstructing the fractal structure of the Hofstadter butterfly spectrum
It is well known that the problem of Bloch electrons on a 2D lattice subject to a perpendicularly applied magnetic field is described by the Hofstadter model. When the magnetic field reaches a strength that allows one magnetic flux quantum per unit cell to pass through, the corresponding energy spectrum versus the dimensionless magnetic flux form a famous fractal structure known as the Hofstadter butterfly [55]. Nevertheless, a magnetic field of approximately $10^{4}$ is necessary to observe the Hofstadter butterfly in typical crystals, which is beyond the capabilities of current experimental techniques. One potential solution to this challenge is the use of superlattices with a larger lattice constant, which could reduce the requisite strength of magnetic fields [70, 71, 72]. On the other hand, the 2D Hofstadter model can be exactly mapped to the 1D diagonal AAH model. Consequently, the Hofstadter butterfly spectrum has also been successfully simulated in 1D AAH-type chains based on superconducting circuits [62, 61]. In this approach, the modulation frequency $\beta$ is used to analogize the required magnetic flux, thereby eliminating the need for a complex 2D lattice and an external magnetic field.
@@
As previously demonstrated in Sec. III.2, in the AAH-type qubit array based on giant-atom wQED, the majority of the resonance dips (corresponding to the energy levels of the AAH model) in the transmission spectrum can be well resolved. Therefore, the transmission spectrum of this system can be effectively utilized to reconstruct the complex and fine fractal structure of the Hofstadter butterfly spectrum with a high degree of fidelity, as shown in Fig. 4(a). In this panel, the number of giant atoms is set to $N=30$. The reduced magnetic flux $\beta$ is varied from 00 to $1$, generating $299$ instances of transmission spectra, which exhibit clear fractal features and a shape that is reminiscent of a butterfly. Figure 4(b) depicts the Hofstadter butterfly energy spectrum obtained from the theoretical model, which exhibits a high degree of agreement with the result in Fig. 4(a). Thus, it appears that an alternative or potentially more efficient approach to simulating the Hofstadter butterfly spectrum is feasible.
@@ III.4 Probing the localization transition
The localization transition [43, 44, 48] is another interesting topic in the incommensurate diagonal AAH model. When the on-site potential of the diagonalized AAH model is made quasiperiodic by setting $\beta$ to an irrational number, the model exhibits a localization transition, where all bulk eigenstates are extended for $V_{0}<2J$, and localized for $V_{0}>2J$. Here we will present an intriguing correlation between the localization transition of the atomic chain and the photon transport behavior observed in the waveguide.
@@
To analyze the relationship between the localization transition and the scattering spectrum, we let the value of the modulation frequency be irrational, with $\beta=(\sqrt{5}-1)/2$, and present in Fig. 5(a) the square of the modulus of the wave functions $\psi_{m}^{(n)}$ (where $m$ and $n$ are used to index the sites and eigenstates, respectively) of the diagonal AAH model for $V_{0}=J/2$, which corresponds to the extended phase. The resulting transmission spectrum and the corresponding Lorentzian decompositions are illustrated in Fig. 5(b). These collective modes with extended feature comprise a small number of superradiant states (with $\tilde{\Gamma}_{n}>\Gamma_{\text{eff}}$) and a significant proportion of subradiant states (with $\tilde{\Gamma}_{n}<\Gamma_{\text{eff}}$), exhibiting a wide range of linewidths, as illustrated in the lower panel of Fig. 5(b). While for $V_{0}=8J$, the situation is quite different. As shown in Fig. 5(c), the wave functions of the atomic chain are highly localized. In this case, the decay of each collective mode to the waveguide is primarily attributable to the most excited atom, which has an excitation probability close to one. Therefore, the width of each resonance dip in the transmission spectrum [see the upper panel of Fig. 5(d)] and the width of each collective excitation amplitude [see the lower panel of Fig. 5(d)] are approximately equal to the effective decay rate $\Gamma_{\text{eff}}$ of a single atom. Namely, all collective states exhibit neither a markedly superradiant nor a distinctly subradiant character.
@@
A more straightforward illustration of the decay rates of the collective modes as a function of index $n$ in the extended and localized phases can be found in Figs. 5(e) and 5(f), respectively. It can be observed that there exists an interesting correspondence between the extended (localized) phase of the atom array and the “localization” (“delocalization”) of the decay rates of its collective modes.
We now seek ways to quantitatively characterize this correspondence. Note that the degree of localization of the $n$th eigenstate of the AAH chain can be described by the inverse participation ratio (IPR):
$\text{IPR}(n)=\sum_{m}\left|\psi^{(n)}_{m}\right|^{4},$ (18)
@@
which vanishes for extended states [see the curves marked with squares in Figs. 6(a) and 6(b)]. As for the aspect of the collective modes, the mean value of their decay rates is equal to the decay rate of a single atom, with $(\sum_{n}\tilde{\Gamma}_{n})/N=\Gamma_{\text{eff}}$. It is therefore reasonable to conclude that the variance
$\sigma^{2}=\frac{1}{N}\sum_{n}\left(\tilde{\Gamma}_{n}-\Gamma_{\text{eff}}\right)^{2}$ (19)
@@
may be employed as a means of measuring the concentration of the decay-rate distribution (vs index $n$). It is evident that the variance $\sigma^{2}\simeq 0$ when $V_{0}>2J$, and $\sigma^{2}>0$ when $V_{0}<2J$ [see the curves marked with circles in Fig. 6(a) and 6(b)], indicating the localized and extended phases, respectively. Note that when $V_{0}\to 0$, the order of magnitude of $\sigma^{2}$ is $N\Gamma_{\text{eff}}^{2}$ for an array with $N\gg 1$. This is due to the fact that in this regime, the decay rates of the superradiant modes, which are a few in number and of the order of $N\Gamma_{\text{eff}}$, contribute mainly to the variance. One can further define a “normalized” decay rate $\tilde{\Gamma}_{\text{norm},n}=\tilde{\Gamma}_{n}/(N\Gamma_{\text{eff}})$, with $\sum_{n}\tilde{\Gamma}_{\text{norm},n}=1$ (which is analogous to the normalization condition $\sum_{m}|\psi^{(n)}_{m}|^{2}=1$ of the wave functions). Therefore, it is possible to emulate the IPR for the wave functions [see Eq. (18)] to define the following IPR describing the distribution of the decay rates
$\text{IPR}_{\text{decay}}=\sum_{n}\tilde{\Gamma}^{2}_{\text{norm},n}.$ (20)
@@
The inverse of this quantity indicates the number of the collective modes that can decay appreciably into the waveguide. As shown in Fig. 6, both $\sigma^{2}$ and $\text{IPR}_{\text{decay}}$ can be used to quantify the localization transition. Moreover, in the thermodynamic limit $N\to\infty$, we have $\sigma^{2}=N\Gamma^{2}_{\text{eff}}\times\text{IPR}_{\text{decay}}$. This is consistent with the results (see the curves marked with circles and triangles) in Fig. 6(b), where a relatively large value of $N=300$ is chosen. Besides, a comparison of Figs. 6(a) and 6(b) illustrates that a finite-size system comprising a greater number of sites is more effective at exhibiting characteristics associated with quantum phase transitions.